GV is a small accounting firm supporting wealthy individuals in their preparation of annual income tax...

  1. GV is a small accounting firm supporting wealthy individuals in their preparation of annual income tax statements. Every December, GV sends out a short survey to their customers, asking for the information required for preparing the tax statements. Based on 50 years of experience, GV categorizes their cases into the following two groups:

    • Group 1 (new customers): 20 percent of cases
    • Group 2 (repeat customers): 80 percent of cases

    This year, there are 50 income tax statements arriving each week. In order to prepare the income tax statement, GV has three resources. The activities are carried out by the following three persons:

    • An administrative support person that processes every tax statement.
    • Senior accountant (who is also the owner): who processes only tax statements for new customers
    • Junior accountant: who only processes tax statements for repeat customers.

    The administrative support person can process 40 statements per week. The senior accountant can process 10 statements per week. The junior accountant can process 20 statements per week; these statements can be completed more quickly given past information on file about the clients.
  1. Draw the demand matrix for this process
  1. What is the total demand rate for the administrator, senior accountant, and the junior accountant?
  1. Which of the three persons is the bottleneck? Show me the calculation (e.g., Implied Utilization)
  1. What is the flow rate for new customers?
  1. What is the flow rate for repeat customers?
Answer #1

a. What is the total demand rate for the administrator, senior accountant, and the junior accountant?

Answer :- To find bottleneck resource, we need to determine the utilization of each resource

Processing capacity of Administrator = Available time per week / Processing time = 40*60/20 = 120 cases / week

There are 50 cases for each week. So use of Administrator = 50/120 = 41.7%

Handling limit of Senior bookkeeper = 40*60/40 = 60 cases for each week.

New cases are 20% of 50 cases for each week Therefore use of Senior bookkeeper = 20%*50/60 = 16.7%

Handling limit of Junior bookkeeper = 40*60/15 = 160 cases for each week.

Rehash cases are 80% of 50 cases for each week Therefore use of Junior bookkeeper = 80%*50/160 = 25%

Use of Administrator is greatest. Consequently it is the bottleneck asset for 20:80 blend of New and Repeat cases.

b.Which of the three persons is the bottleneck? Show me the calculation (e.g., Implied Utilization)

  1. Administrator              
  2.    Senior accountant              
  3.    Junior accountant  

c. What is the flow rate for new customers?

Answer :- Flow rate of New cases per week = 50*20% = 10 cases / week

d. What is the flow rate for repeat customers?

Answer :-  Flow rate of Repeat cases per week = 50*80% = 40 cases / week